OpenStudy (anonymous):

the solution to the equation 2x + 3 = 3x + ? is x = -8. What is the number on the question mark?

OpenStudy (solomonzelman):

plug in -8 for x, can you do that?

OpenStudy (smored):

so let's make ?=y. 2(8)+3=3(8)+y 16+3=24+y y=-5

OpenStudy (solomonzelman):

@smored, you are plugging -8, not 8.

OpenStudy (smored):

oops

OpenStudy (solomonzelman):

don't do his work please.

OpenStudy (smored):

so -16+3=-24+y y=11

OpenStudy (smored):

sorry for my mistake